# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v3

@@ F.2 Positive Noise on TinyImageNet
Table 13: Comparison with SOTA methods on Office-Home. The best performance is marked in red.
@@
ResNet-50[22] 44.9 66.3 74.3 51.8 61.9 63.6 52.4 39.1 71.2 63.8 45.9 77.2 59.4
MinEnt[21] 51.0 71.9 77.1 61.2 69.1 70.1 59.3 48.7 77.0 70.4 53.0 81.0 65.8
SAFN[65] 52.0 71.7 76.3 64.2 69.9 71.9 63.7 51.4 77.1 70.9 57.1 81.5 67.3
CDAN+E[34] 54.6 74.1 78.1 63.0 72.2 74.1 61.6 52.3 79.1 72.3 57.3 82.8 68.5
DCAN[29] 54.5 75.7 81.2 67.4 74.0 76.3 67.4 52.7 80.6 74.1 59.1 83.5 70.5
BNM [10] 56.7 77.5 81.0 67.3 76.3 77.1 65.3 55.1 82.0 73.6 57.0 84.3 71.1
SHOT[31] 57.1 78.1 81.5 68.0 78.2 78.1 67.4 54.9 82.2 73.3 58.8 84.3 71.8
ATDOC-NA[32] 58.3 78.8 82.3 69.4 78.2 78.2 67.1 56.0 82.7 72.0 58.2 85.5 72.2
ViT-B[15] 54.7 83.0 87.2 77.3 83.4 85.6 74.4 50.9 87.2 79.6 54.8 88.8 75.5
TVT-B[67] 74.9 86.8 89.5 82.8 88.0 88.3 79.8 71.9 90.1 85.5 74.6 90.6 83.6
CDTrans-B[66] 68.8 85.0 86.9 81.5 87.1 87.3 79.6 63.3 88.2 82.0 66.0 90.6 80.5
SSRT-B [57] 75.2 89.0 91.1 85.1 88.3 90.0 85.0 74.2 91.3 85.7 78.6 91.8 85.4
@@
ResNet-50[22] 55.1 53.3 61.9 59.1 80.6 17.9 79.7 31.2 81.0 26.5 73.5 8.5 52.4
DANN[19] 81.9 77.7 82.8 44.3 81.2 29.5 65.1 28.6 51.9 54.6 82.8 7.8 57.4
MinEnt[21] 80.3 75.5 75.8 48.3 77.9 27.3 69.7 40.2 46.5 46.6 79.3 16.0 57.0
SAFN[65] 93.6 61.3 84.1 70.6 94.1 79.0 91.8 79.6 89.9 55.6 89.0 24.4 76.1
CDAN+E[34] 85.2 66.9 83.0 50.8 84.2 74.9 88.1 74.5 83.4 76.0 81.9 38.0 73.9
BNM [10] 89.6 61.5 76.9 55.0 89.3 69.1 81.3 65.5 90.0 47.3 89.1 30.1 70.4
CGDM[16] 93.7 82.7 73.2 68.4 92.9 94.5 88.7 82.1 93.4 82.5 86.8 49.2 82.3
SHOT[31] 94.3 88.5 80.1 57.3 93.1 93.1 80.7 80.3 91.5 89.1 86.3 58.2 82.9
ViT-B[15] 97.7 48.1 86.6 61.6 78.1 63.4 94.7 10.3 87.7 47.7 94.4 35.5 67.1
TVT-B[67] 92.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.9
CDTrans-B[66] 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
SSRT-B [57] 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
@@
F.3 Positive Noise for Domain Adaptation
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions [42] [63]. There are mainly two kinds of deep neural networks for UDA, which are CNN-based and Transformer-based methods [57] [67]. Various techniques for UDA are adopted on these backbone architectures. For example, the discrepancy techniques measure the distribution divergence between source and target domains [34] [55]. Adversarial adaptation discriminates domain-invariant and domain-specific representations by playing an adversarial game between the feature extractor and a domain discriminator [19].
Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [62] and VisDA2017 [43]. Office-Home[62] has 15,500 images of 65 classes from four domains: Artistic (Ar), Clip Art (Cl), Product (Pr), and Real-world (Rw) images. VisDA2017 is a Synthetic-to-Real object recognition dataset, with more than 0.2 million images in 12 classes. We use the ViT-B with a $16\times 16$ patch size, pre-trained on ImageNet. We use minibatch Stochastic Gradient Descent (SGD) optimizer [46] with a momentum of 0.9 as the optimizer. The batch size is set to 32. We initialized the learning rate as 0 and linearly warm up to 0.05 after 500 training steps. The results are shown in Table 13 and 14. The methods above the black line are based on CNN architecture, while those under the black line are developed from the Transformer architecture. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results show that positive noise can improve model generality and, therefore, benefit deep models in domain adaptation tasks.

### v4

@@ F.10 Positive Noise for Domain Adaptation
Table 21: Comparison with SOTA methods (ResNet-50 He et al. (2016), MinEnt Grandvalet & Bengio (2004), SAFN Xu et al. (2019), CDAN+E Long et al. (2018), DCAN Li et al. (2020), BNM Cui et al. (2020), SHOT Liang et al. (2020), ATDOC-NA Liang et al. (2021), ViT-B Dosovitskiy et al. (2020), TVT-B Yang et al. (2023a), CDTrans-B Xu et al. (2022), SSRT-B Sun et al. (2022)) on Office-Home. The best performance is marked in red.
@@
ResNet-50 44.9 66.3 74.3 51.8 61.9 63.6 52.4 39.1 71.2 63.8 45.9 77.2 59.4
MinEnt 51.0 71.9 77.1 61.2 69.1 70.1 59.3 48.7 77.0 70.4 53.0 81.0 65.8
SAFN 52.0 71.7 76.3 64.2 69.9 71.9 63.7 51.4 77.1 70.9 57.1 81.5 67.3
CDAN+E 54.6 74.1 78.1 63.0 72.2 74.1 61.6 52.3 79.1 72.3 57.3 82.8 68.5
DCAN 54.5 75.7 81.2 67.4 74.0 76.3 67.4 52.7 80.6 74.1 59.1 83.5 70.5
BNM 56.7 77.5 81.0 67.3 76.3 77.1 65.3 55.1 82.0 73.6 57.0 84.3 71.1
SHOT 57.1 78.1 81.5 68.0 78.2 78.1 67.4 54.9 82.2 73.3 58.8 84.3 71.8
ATDOC-NA 58.3 78.8 82.3 69.4 78.2 78.2 67.1 56.0 82.7 72.0 58.2 85.5 72.2
ViT-B 54.7 83.0 87.2 77.3 83.4 85.6 74.4 50.9 87.2 79.6 54.8 88.8 75.5
TVT-B 74.9 86.8 89.5 82.8 88.0 88.3 79.8 71.9 90.1 85.5 74.6 90.6 83.6
CDTrans-B 68.8 85.0 86.9 81.5 87.1 87.3 79.6 63.3 88.2 82.0 66.0 90.6 80.5
SSRT-B 75.2 89.0 91.1 85.1 88.3 90.0 85.0 74.2 91.3 85.7 78.6 91.8 85.4
@@
ResNet-50He et al. (2016) 55.1 53.3 61.9 59.1 80.6 17.9 79.7 31.2 81.0 26.5 73.5 8.5 52.4
DANNGanin & Lempitsky (2015) 81.9 77.7 82.8 44.3 81.2 29.5 65.1 28.6 51.9 54.6 82.8 7.8 57.4
MinEntGrandvalet & Bengio (2004) 80.3 75.5 75.8 48.3 77.9 27.3 69.7 40.2 46.5 46.6 79.3 16.0 57.0
SAFNXu et al. (2019) 93.6 61.3 84.1 70.6 94.1 79.0 91.8 79.6 89.9 55.6 89.0 24.4 76.1
CDAN+ELong et al. (2018) 85.2 66.9 83.0 50.8 84.2 74.9 88.1 74.5 83.4 76.0 81.9 38.0 73.9
BNM Cui et al. (2020) 89.6 61.5 76.9 55.0 89.3 69.1 81.3 65.5 90.0 47.3 89.1 30.1 70.4
CGDMDu et al. (2021) 93.7 82.7 73.2 68.4 92.9 94.5 88.7 82.1 93.4 82.5 86.8 49.2 82.3
SHOTLiang et al. (2020) 94.3 88.5 80.1 57.3 93.1 93.1 80.7 80.3 91.5 89.1 86.3 58.2 82.9
ViT-BDosovitskiy et al. (2020) 97.7 48.1 86.6 61.6 78.1 63.4 94.7 10.3 87.7 47.7 94.4 35.5 67.1
TVT-BYang et al. (2023a) 92.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.9
CDTrans-BXu et al. (2022) 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
SSRT-B Sun et al. (2022) 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8